shift summary: owl shift 10/14/99

John Shields (jshields@ksera)
Thu, 14 Oct 1999 07:02:57 -0500

Shift Summary for owl shift, 10/14/99

On Shift: J. Shields, Hogan

Hours of beam available at all intensities: 6.5/7.0
Hours of beam available at nominal intensity: 6.4
Hours of E799 data (special runs not included): 6.5
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:

tevatron down for first 30 min of shift
========================================================================
Runs ended this shift:
Run 14834 - ( lousy data) - BEAM_799

Tapes brought to FCC this shift:
(including tapes for online split)

none
========================================================================
Accesses during shift:

none

========================================================================
Overview of shift:
We got off to a rocky start by facing busy triggers, dead processes, and
unhappy TRDs all simultaniously. However, once we finally cleared these
hurdles, datataking became relatively smooth.
Unfortunately, these problems probably make the first 2 hours of data
more or less worthless.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

14834 00:30 beam returns and a new run begins

00:31 The fun begins! Repeated busys from banana 2X.
kfrending and resetting bananas have no
discerable impact. Jason L. paged

00;45 Jason L. dutifully arrives to debug bananas.
Meanwhile the intrepid shift crew valiantly battle
SDAQ red buttons (for dead processes trgchk and
scaler_check) and TRD alarms. Pandimonium reigns
as shiftmates literally fight over phones and
keyboards. (Well, ok fine. They didn't *literally*
fight or anything exciting like that, but you
get the idea...)

02:45 after extensive debugging and then noticing that
an entire quadrant of the csi was missing from the
monitoring, crew decides to officially write off
14834 as a worthless run and to start anew.

14835 02:50 new run begins. As suspected, CSI quadrant
reappears on the monitoring display.

03:05 JL reaches a verdict as to why the bananas were
unhappy: it appears that the rack door was pinching
the cable, so the door is now temporarily being left
open. DO NOT SHUT DOOR RR177!

03:55 CSI # 947 has large RMS (30%). Ok a few spills later.

06:40 on-split request. UPK009 mounted on ktev1.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Page Jason L. if there is a downtime longer than an hour.